# A Quantitative Comparison of Pre-Trained Model Registries to Traditional Software Package Registries

### v1

@@ Abstract.
Background: Software Package Registries (SPRs) are an integral part of the software supply chain. These collaborative platforms unite contributors, users, and packages, and they streamline package management. Much engineering work focuses on synthesizing packages from SPRs into a downstream project. Prior work has thoroughly characterized the SPRs associated with traditional software, such as NPM (JavaScript) and PyPI (Python). Pre-Trained Model (PTM) Registries are an emerging class of SPR of increasing importance, because they support the deep learning supply chain. Aims: A growing body of empirical research has examined PTM registries from various angles, such as vulnerabilities, reuse processes, and evolution. However, no existing research synthesizes them to provide a systematic understanding of the current knowledge. Furthermore, much of the existing research includes unsupported qualitative claims and lacks sufficient quantitative analysis. Our research aims to fill these gaps by providing a thorough knowledge synthesis and use it to inform further quantitative analysis.
Methods: To consolidate existing knowledge on PTM reuse, we first conduct a systematic literature review (SLR). We then observe that some of the claims are qualitative and lack quantitative evidence. We identify quantifiable metrics assoiated with those claims, and measure in order to substantiate these claims. Results: From our SLR, we identify 12 claims about PTM reuse on the HuggingFace platform, 4 of which lack quantitative validation. We successfully test 3 of these claims through a quantitative analysis, and directly compare one with traditional software. Our findings corroborate qualitative claims with quantitative measurements. Our two most notable findings are: (1) PTMs have a significantly higher turnover rate than traditional software, indicating a dynamic and rapidly evolving reuse environment within the PTM ecosystem; and (2) There is a strong correlation between documentation quality and PTM popularity. Conclusions: Our findings validate several qualitative research claims with concrete metrics, confirming prior qualitative and case study research. Our measures show further dynamics of PTM reuse, motivating further research infrastructure and new kinds of measurements.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
As the size and cost of developing deep learning (DL) models from scratch continue to rise, engineers are increasingly turning to adapt open-source Pre-trained Models (PTMs) as a cost-effective alternative (Jiang et al., 2022). PTM registries facilitate the reuse of open-source models by providing packages that include pre-trained weights, configuration, and documentation (Jiang et al., 2023b). Hugging Face has become a prominent PTM registry, comparable in popularity to traditional software registries like NPM and PyPI (Jiang et al., 2023b).
Prior research has empirically compared PTM registries to traditional software package registries such as NPM and PyPI, tackling diverse issues such as carbon emissions, model selection, and vulnerabilities (Jiang et al., 2023b; Castaño et al., 2023; Kathikar et al., 2023). Despite these efforts, no work has yet synthesized the existing knowledge of PTM registries. Furthermore, not all topics about PTM registries such as Hugging Face have been studied from both qualitative and quantitative perspectives. This lack of a comprehensive approach has led to a lack of quantitative validation, leaving numerous qualitative insights about PTM registries under-verified.
@@
Figure 1. Overview of this work’s context and approach. Much is known, both qualitatively and quantitatively, about package reuse processes in traditional software package registries such as NPM and PyPI. Meanwhile, empirical data about the reuse of pre-trained deep neural network models (PTMs) is emerging. This work provides the first systematic literature review on PTM reuse, focused on extracting the claims present in the prior work (RQ1) and providing quantitative evaluation of the un-quantified and under-quantified claims (RQ2).
Our work aims to bridge the knowledge gap concerning the synthesis of extensive research on PTM reuse facilitated by Hugging Face, and provide further measurements to validate the prior qualitative and quantitative claims. An overview is shown in Figure 1. Our work consists of two parts: First, we conduct a systematic literature review to transform qualitative insights about PTM package reuse on Hugging Face into quantitative metrics. Second, we evaluate the robustness of these insights through quantitative validation. This approach enhances our understanding of package reuse dynamics on Hugging Face compared to traditional SPRs.
@@
Our findings indicate that most of the claims from prior qualitative results are correct. Of the 3 prior claims we quantitatively evaluated, 2 were supported by our results, and 1 was supported by 1 measurement and not supported by the other. Our measurement of library usage indicated a preference for using the Transformers library when creating descendents of models, with 80% of descendents of models choosing to utilize Transformers. Our measurement of package turnover indicates that HuggingFace has a significantly higher turnover rate than traditional package registries. Our measurement of model popularity and descendent count found a correlation between the two, further supporting claims that popularity is a driver of model selection. Finally, we found a strong correlation between documentation quality and PTM popularity, with the top 1000 models outperforming the bottom 1000 in documentation.
@@
  • We conduct a systematic literature review on PTM reuse in the Hugging Face registry, and extract a list of qualitative claims from prior work. (section 4)
  • We map the qualitative claims to quantitative measurements. We use these measurements to validate the prior findings, via comparison of our quantitative measurements of Hugging Face to representative traditional SPRs. (section 5)
  • Our work provides recommendations for future work on developing tools to analyze and keeping datasets updated to support further investigations on the PTM supply chain. (section 7)
Significance for Software Engineering: Prior work has made both quantitative and qualitative claims about software engineers’ reuse of PTMs. We systematically synthesized this knowledge through a literature review and developed quantifiable metrics to corroborate the qualitative claims. Our findings confirm several qualitative results about PTM reuse, and also quantify the dynamic reuse environment within the PTM ecosystem. Our results will inform research infrastructure and the development of new metrics to guide and refine PTM development and reuse.
@@ 2. Background and Related Work
In this section, we first explain how software package registries facilitate the reuse of software (section 2.1). We then discuss the rise of Deep Neural Networks (DNNs) and their presence in software engineering reuse as PTM packages (section 2.2). We also detail how software package registries have been measured in previous works, and the limited effort towards quantitatively evaluating PTM registries (section 2.3). Our work aims to advance the state of the art when it comes to quantitative evaluation of PTM registries.
2.1. Software Package Registries
Figure 2. The Software Supply Chain. Engineers contribute Packages to model registries, Packages evolve through internal development and external dependence on other Packages, and are used by downstream Reusers who incorporate them into Systems.
Figure 2 depicts the software supply chain. Software Package Registries (SPRs) serve as collaborative hubs that connect package contributors, reusers, and the packages themselves, facilitating software reuse. These registries are important for engineers because they provide comprehensive information that significantly enhances downstream software development.
SPRs act as platforms for creators to upload and share their software packages, featuring version control to allow users to access previous package versions as needed (npm, [n. d.]; pyp, [n. d.]). These platforms promote package discoverability and perceived quality through user engagement tools like comments, likes, and ratings. Additionally, collaboration is fostered by discussion threads and version management, enhancing user interaction and visibility. Registries also provide comprehensive metadata, improving package visibility in search results and aiding in package selection. Tools for package downloading, bundling, and managing versions and updates streamline the reuse process, simplifying package lifecycle management.
Recently, PTM package registries (e.g., Hugging Face (Wolf et al., 2020), PyTorch Hub (pyt, [n. d.]), ONNX Model Zoo (onn, [n. d.])) emerged to support efficient development of AI systems (Davis et al., 2023; Jiang et al., 2022). PTM packages include traditional components (e.g., documentation, dependencies). However, they also include additional DNN-specific components, such as pre-trained weights, training dataset, and model architecture (Jiang et al., 2023b). Figure 3 depicts the structural similarity in package reuse between traditional and PTM packages.
Figure 3. Reuse processes of traditional software and of PTMs, as reported by Jiang et al. (Jiang et al., 2023b). Reuse processes are similar, suggesting that SPR measurements and trends may be similar.
2.2. Deep Neural Networks and Pre-Trained Model Packages
@@
Deep Neural Networks (DNNs), comprising numerous hidden layers, have gained popularity as a cutting-edge solution for complex problems in various fields, such as image recognition in autonomous vehicles (Garcia et al., 2020) and AI voice assistant systems (Nasirian et al., 2017). Developing and training these models from scratch requires significant time and resources (Rae et al., 2021; Hoffmann et al., 2022; Patterson, 2022). For example, a Llama-2-70B model needs 1720K GPU hours to train from scratch (Touvron et al., 2023). To overcome this, engineers increasingly opt for using Pre-Trained Models (PTMs), which allows them to forego the lengthy and resource-intensive initial training phase. By leveraging PTMs, they can focus on a much shorter training period to fine-tune the models for specific tasks (Jiang et al., 2023b; Davis et al., 2023; Jiang et al., 2023a).
@@
Within the context of PTMs, the term “software reuse” refers to the reuse of models along with their training configurations and weights, i.e., reuse of the PTM packages. Figure 3 shows the comparison of reuse process between PTM and traditional packages. Reusing DNNs as PTM packages mirrors traditional practices of software package reuse, where existing software components are integrated and adapted rather than built from scratch (Jiang et al., 2022; Gong et al., 2023a). This practice can enhance efficiency and reducing development time (Gopalakrishna et al., 2023). This adaptation involves not only the models but also their pre-trained states, which can be adjusted for specific applications, thereby constituting a combination of model reuse and customization rather than traditional software package reuse alone.
2.3. Measurements of Software Package Registries
Research on software package registries has traditionally focused on quantifying various aspects (Zerouali et al., 2019a; Decan et al., 2018; German et al., 2010). However, PTM registry research has primarily focused on qualitative and small-scale quantitative measurements.
2.3.1. General Comparisons:
@@
This research approach has been adapted to PTM registries recently (Jiang et al., 2024a, 2023b). Early studies have explored how PTMs are reused and adapted, examining both qualitative and quantitative aspects. For instance, research has started to analyze contribution patterns and the reuse dynamics specific to PTM registries, such as those found in TensorFlow Hub and Hugging Face (Jiang et al., 2023b; Miao et al., 2017; Davis et al., 2023). However, there are still gaps in understanding the evolution and reuse patterns in the PTM registries (Jiang et al., 2023b, 2024b). Our work bridges this gap by providing additional quantitative measurements and comparing our results to traditional SPRs.
2.3.2. To Explain or Quantify Phenomena:
Despite these efforts, there remains a gap in research that connects qualitative observations with quantitative measurements across these registries.
Recent studies within traditional SPRs typically start with qualitative observations that are later supported by quantitative data. Issues such as package obsolescence and the spread of vulnerabilities, initially observed quantitatively, have been rigorously quantified in ecosystems like NPM (Mujahid et al., 2023; Abdalkareem et al., 2017). These investigations form a crucial basis for comprehending how features of software registries impact the practices of software development.
In the research domain of PTM registries, similar investigations are needed. For example, Jiang et al. provide a comprehensive analysis on the risks while using PTMs, and the PTM reuse process (Jiang et al., 2022, 2023b). As a follow-up work, they also collected both qualitative and quantitative data on PTM naming practices (Jiang et al., 2024a). This line of research is crucial for understanding PTM package registries. However, their qualitative insights have not been substantiated by quantitative analysis. We address this gap by conducting a systematic literature review and deriving quantifiable metrics to assess the claims from previous studies.
@@ 3. Knowledge Gap and Research Questions
To summarize the knowledge gap, we lack a cohesive understanding on PTM package reuse that synthesizes prior research on Hugging Face. Additionally, there is a lack of some quantitative measurements, which reduce our understanding of the registry.
To address the gap of knowledge synthesis, we ask:
@@
By gathering and analyzing these claims, we aim to convert the qualitative data into a set of quantifiable metrics. This effort will not only enrich our understanding of Hugging Face as a PTM platform but also enhance the comparability of data across different software ecosystems. Specifically, this analysis enables future researchers to establish metrics that facilitate the comparison of Hugging Face and traditional software registries similarities and differences in package reuse practices.
@@ 3.1. Overview of Methodology
Figure 1 shows the overview of our work’s context and approach. We answered these questions through three steps:
@@
    We conducted a systematic literature review (SLR) on PTM reuse, aiming to compile a comprehensive list of claims about package reuse.
@@
    We synthesized these claims into quantifiable measurements, performed the measurements, and then compared them to the previous findings.
@@
    Where possible, we compared these to those from traditional software to see differences between the reuse patterns in different registries.
@@
4. RQ1: What claims about package reuse on Hugging Face are made by prior research?
@@
To enhance our understanding of the existing claims related to PTM reuse within package registries, we followed empirical standards and conducted a systematic literature review (Ralph et al., 2020; Ralph and Baltes, 2022; Borenstien et al., 2009; Keele et al., 2007). Our initial step involved a pilot study to define the scope of our review (section 4.1.1). A systematic literature review entails five steps: identification of research, study selection, study quality assessment, data extraction, and data analysis (Keele et al., 2007; Moher et al., 2010). Figure 4 illustrates the results of each step. We detail them next.
Figure 4. Systematic literature review process. Step 1 details the search query used. The subsequent filtering and distinct claims, along with the summarized categories, are discussed in section 4.2.

### v2

# What do we know about Hugging Face? A systematic literature review and quantitative validation of qualitative claims

@@ Abstract.
Background: Software Package Registries (SPRs) are an integral part of the software supply chain. These collaborative platforms unite contributors, users, and code for streamlined package management. Prior work has characterized the SPRs associated with traditional software, such as NPM (JavaScript) and PyPI (Python). Pre-Trained Model (PTM) Registries are an emerging class of SPR of increasing importance, because they support the deep learning supply chain. A growing body of empirical research has examined PTM registries from various angles, such as vulnerabilities, reuse processes, and evolution. However, no synthesis provides a systematic understanding of current knowledge. Furthermore, much of the existing research includes non-quantified qualitative observations.
Aims: First, we aim to provide a systematic knowledge synthesis. Second, we quantify qualitative claims.
Methods: We conducted a systematic literature review (SLR). We then observed that some of the claims are qualitative, lacking quantitative evidence. We identify quantifiable metrics associated with those claims, and measure in order to substantiate these claims.
Results: We identify 12 claims about PTM reuse on the HuggingFace platform, 4 of which lack quantitative support. We tested 3 of these claims through a quantitative analysis, and directly compare the fourth with traditional software. Our most notable findings are: (1) PTMs have a significantly higher turnover rate than traditional software, indicating more rapid evolution; and (2) There is a strong correlation between documentation quality and PTM popularity.
Conclusions: Our findings validate several qualitative research claims with concrete metrics, confirming prior research. Our measures motivate further research on the dynamics of PTM reuse.
Software Package Registries, Software Supply Chain, Pre-Trained Models, Empirical Software Engineering, Literature Review
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 18th ACM / IEEE International Symposium on Empirical Software Engineering and Measurement; October 24–25, 2024; Barcelona, Spainbooktitle: Proceedings of the 18th ACM / IEEE International Symposium on Empirical Software Engineering and Measurement (ESEM ’24), October 24–25, 2024, Barcelona, Spaindoi: 10.1145/3674805.3686665isbn: 979-8-4007-1047-6/24/10ccs: Software and its engineering Software libraries and repositoriesccs: General and reference Surveys and overviews
@@ 1. Introduction
Figure 1. Overview of this work’s context and approach. Pre-trained deep neural network models (PTMs) are beginning to be reused, and accompanying empirical research is emerging. This work provides the first systematic literature review on PTM reuse. We extract the claims in prior work (RQ1) and provide quantitative evaluation of un-quantified and under-quantified claims (RQ2).
As the size and cost of developing deep learning (DL) models from scratch continue to rise, engineers are increasingly turning to adapt open-source Pre-trained Models (PTMs) as a cost-effective alternative (Jiang et al., 2022). PTM registries facilitate the reuse of open-source models by providing packages that include pre-trained weights, configuration, and documentation (Jiang et al., 2023a). Hugging Face has become a prominent PTM registry, comparable in popularity to traditional software registries like NPM and PyPI (Jiang et al., 2023a). Understanding the characteristics of PTM registries such as Hugging Face is key to supporting effective and efficient software reuse in this emerging context.
Prior research has made significant strides in comparing PTM registries to traditional software package registries, addressing issues like carbon emissions, model selection, and vulnerabilities (Jiang et al., 2023a; Castaño et al., 2024; Kathikar et al., 2023). However, there has been no systematic literature review that describes the current state of knowledge; such reviews advance the field by providing research agendas. Our study thus contributes in three ways. First, we provide the first systematic review of knowledge on PTM registries. Second, we propose quantitative metrics for existing qualitative insights, enabling a more robust validation of existing claims about PTM registries. Lastly, we validate or challenge previous qualitative insights through quantitative analysis.
As illustrated in Figure 1, our method has two parts. First, we conduct a systematic literature review (SLR) to extract existing knowledge (claims) about Hugging Face. Second, we identify un-quantified and under-quantified claims, and provide metrics and measurements using existing datasets.
@@
Our SLR extracted 12 distinct claims about HuggingFace, of which 4 lacked large-scale quantitative evidence. After defining metrics, we supported 2 of these with a large-scale measure; and found ambivalent evidence for a third. The fourth was too imprecise to operationalize. To summarize our findings: (1) Our measurement of library usage indicated a preference for using the Transformers library when creating descendents of models, with 80% of descendents of models choosing to utilize Transformers ; (2) HuggingFace has a significantly higher turnover rate than traditional package registries; (3) Model popularity correlates with descendent count, supporting claims that popularity drives model selection; and (4) Documentation quality correlates with PTM popularity.
@@
  • We conduct a systematic literature review on PTM reuse in Hugging Face, summarizing the claims of prior work (section 4).
  • We operationalize qualitative claims with quantitative measurements, and then validate most prior qualitative findings, via comparison of our quantitative measurements of Hugging Face to representative traditional SPRs (section 5).
  • We recommend future work on analyzing and maintaining datasets for furrther study of the PTM supply chain (section 7).
Significance: Prior work has made both quantitative and qualitative claims about software engineers’ reuse of PTMs. We systematically synthesized this knowledge through a literature review, and developed quantifiable metrics to corroborate qualitative claims. Our findings confirm several qualitative results about PTM reuse, and also quantify the dynamic reuse environment within the PTM ecosystem. Our results will inform research infrastructure, and new metrics to guide and refine PTM development and reuse.
@@ 2. Background and Related Work
We discuss how software package registries facilitate the reuse of software, and deep neural network-specific considerations for package registries that contain Pre-Trained Models (section 2.1). We describe how software package registries have been measured in previous works, and the limited quantification of PTM registries (section 2.2).
2.1. Software Package Registries, Old and New
Figure 2. The software supply chain. Software package registries connect package authors to reusers, accelerating system development.
2.1.1. General Concepts.
Software Package Registries (SPRs) serve as collaborative hubs that connect package contributors to package reusers, facilitating software reuse. Figure 2 depicts how contributors collaborate on packages that are reused as components further down the software supply chain. For package contributors (maintainers), SPRs provide tools for package lifecycle management: bundling, versioning, hosting, and so on (npm, [n. d.]; pyp, [n. d.]). For package reusers, SPRs assist in package search and retrieval, and often provide direct or indirect measures of package quality through user engagement tools like comments, likes, and ratings. SPRs typically provide both existing components, and metadata about their provenance, performance, and other attributes that constrain the potential for reuse.
2.1.2. SPRs for Deep Neural Networks.
@@ 2.1. Software Package Registries, Old and New
Deep Neural Networks (DNNs) provide state-of-the-art performance for many tasks, such as image recognition in autonomous vehicles (Garcia et al., 2020) and AI voice assistant systems (Nasirian et al., 2017). Developing and training these models from scratch requires significant time and resources (Rae et al., 2021; Hoffmann et al., 2022; Patterson, 2022). For example, a Llama-2-70B model needs 1720K GPU hours to train from scratch (Touvron et al., 2023). Engineers thus often use Pre-Trained Models (PTMs). They can then undertake a much shorter training period to fine-tune the models for specific tasks (Jiang et al., 2023a; Davis et al., 2023; Jiang et al., 2024a).
Since PTMs are reused components, PTM package registries (e.g., Hugging Face (Wolf et al., 2020), PyTorch Hub (pyt, [n. d.]), ONNX Model Zoo (onn, [n. d.])) have emerged to support the development and reuse of trained DNNs for use in the development of AI systems (Davis et al., 2023; Jiang et al., 2022). PTM package registries are modeled on traditional SPRs, with some extensions for DNN-specific properties. PTM packages include traditional package components such as documentation and dependencies. However, they also include additional DNN-specific components, such as pre-trained weights, training dataset(s), and model architecture (Jiang et al., 2023a).
@@
Within the context of PTMs, the term “software reuse” refers to the reuse of models along with their training configurations and weights, i.e., reuse of the PTM packages. Figure 3 compares the reuse process between traditional packages and PTM packages. Reusing DNNs as PTM packages mirrors traditional practices of software package reuse, where existing software components are integrated and adapted rather than built from scratch (Jiang et al., 2022; Gong et al., 2023). This practice can enhance efficiency and reducing development time (Gopalakrishna et al., 2023). This adaptation involves not only the models but also their pre-trained states, which can be adjusted for specific applications, thereby constituting a combination of model reuse and customization rather than traditional software package reuse alone.
Figure 3. Reuse processes of traditional software and of PTMs, as reported by Jiang et al. (Jiang et al., 2023a). Reuse processes are similar, suggesting that SPR measurements and trends may be similar.
2.2. Measuring Software Package Registries
Research on traditional software package registries has typically sought to quantify various aspects of SPRs (Zerouali et al., 2019a; Decan et al., 2018; German et al., 2010). In contrast, PTM registry research has often focused on qualitative and small-scale quantitative measurements to date. This difference is appropriate given the need to provide qualitative descriptions of phenomena prior to measurement of them. However, this difference also highlights the novel aspects of PTM registries and emphasizes the need for the systematization of existing knowledge.
2.2.1. General Comparisons.
@@ 2.2. Measuring Software Package Registries
This research approach has been adapted to PTM registries recently (Jiang et al., 2024b, 2023a). Early studies have explored how PTMs are reused and adapted, examining both qualitative and quantitative aspects. For instance, research has started to analyze contribution patterns and the reuse dynamics specific to PTM registries, such as those found in TensorFlow Hub and Hugging Face (Jiang et al., 2023a; Miao et al., 2017; Davis et al., 2023). However, there are still gaps in understanding the evolution and reuse patterns in the PTM registries (Jiang et al., 2023a, 2024c). Our work bridges this gap by providing additional quantitative measurements and comparing our results to traditional SPRs.
2.2.2. Explaining Phenomena Before Quantifying Them.
Despite these efforts, there remains a gap in research that connects qualitative observations with quantitative measurements in PTM registries. Recent studies within traditional SPRs typically start with qualitative observations that are later supported by quantitative data. Issues such as package obsolescence and the spread of vulnerabilities, initially observed qualitatively, have been rigorously quantified in ecosystems such as NPM (Mujahid et al., 2023; Abdalkareem et al., 2017). These investigations provide a basis for understanding how aspects of software registries impact the practices of software development.
In the research domain of PTM registries, similar investigations are needed. For example, Jiang et al. provide a comprehensive analysis on the risks while using PTMs, and the PTM reuse process (Jiang et al., 2022, 2023a). As a follow-up work, they also collected both qualitative and quantitative data on PTM naming practices (Jiang et al., 2024b). This line of research is crucial for understanding PTM package registries. However, their qualitative insights have not been substantiated by quantitative analysis. We address this gap by conducting a systematic literature review and deriving quantifiable metrics to assess the claims from previous studies.
@@ 3. Knowledge Gap and Research Questions
To summarize the knowledge gap, we lack a cohesive understanding on PTM package reuse that synthesizes prior research on Hugging Face. Additionally, there is a lack of some quantitative measurements, which reduce our understanding of the registry. To address the gap of knowledge synthesis, we ask:
@@
By gathering and analyzing these claims, we convert qualitative data into a set of quantifiable metrics. This effort enriches our understanding of Hugging Face as a PTM platform and enhances the comparability of data across different software ecosystems. Specifically, future researchers are enabled to establish metrics that facilitate the comparison of Hugging Face and traditional software registries similarities and differences in package reuse practices.
@@ 3.1. Overview of Methodology
Figure 1 shows the overview of our work’s context and approach.
@@
    We conducted a systematic literature review on PTM reuse, compiling a comprehensive list of claims about package reuse.
@@
    We synthesized these claims into quantifiable metrics, took measurements, and compared them to previous findings.
@@
    We compared measurements of PTM packages to those of traditional packages to see differences in reuse patterns.
@@
4. RQ1: Prior Package Reuse Claims
@@
To enhance our understanding of the existing claims related to PTM reuse within package registries, we conducted a systematic literature review (SLR) following established empirical standards (Ralph et al., 2020; Ralph and Baltes, 2022; Borenstien et al., 2009; Kitchenham and Charters, 2007). The SLR protocol was designed to ensure a comprehensive and unbiased review of the literature, with a focus on identifying quantifiable metrics that could be used to assess claims about PTM reuse on Hugging Face.
The need for quantifiable metrics in this study stems from the importance of objectively measuring and comparing aspects of PTM reuse across different studies and contexts. These metrics allow us to validate qualitative observations with quantitative data, identify trends and patterns in PTM reuse practices, and compare findings across different studies and ecosystems.
Our initial step involved a pilot study to define the scope of our review (section 4.1.1). A systematic literature review entails five steps: identification of research, study selection, study quality assessment, data extraction, and data analysis (Kitchenham and Charters, 2007; Moher et al., 2010). Figure 4 illustrates the results of each step. We detail them next.
Figure 4. Systematic literature review process and results.